if the coordinates of two adjacent corners of a square at at A(4,2) and B(7,6) what is the length of the diagonal?
Answer by Alan3354(21605)

if the coordinates of two adjacent corners of a square at at A(4,2) and B(7,6) what is the length of the diagonal?
--------------------
The length of the side is s.
s^2 = diffy^2 + diffx^2
s^2 = (6-2)^2 + (7-4)^2
s^2 = 4^2 + 3^2 = 16 + 9
s^2 = 25
s = 5
---------
Since it's a square, all sides = 5
Use Pythagoras,
d = 5sqrt(2)